# Dual axis atomic magnetometer and gyroscope enabled by nuclear spin perturbation

### v1

@@ I Introduction
The gyroscopic performance of the SCC has shown excellent levels of sensitivity. For example Kornack et al. [12] found a noise floor $\mathrm{5\!\times\!10^{-7}\,rad\,s^{-1}\,Hz^{-1/2}}$, which was recently surpassed by Wei et al. [23] with $\mathrm{3\!\times\!10^{-8}\,rad\,s^{-1}\,Hz^{-1/2}}$. While impressive, these results are still over an order of magnitude above the atomic shot noise limit set by the alkali, and particularly limited at low frequencies where magnetic noise is typically largest. While some progress has been made reducing magnetic and other technical noise sources [24, 25, 26, 27, 28], a remaining challenge is that increased suppression of magnetic noise is not only hard to achieve, it is also difficult to maintain this suppression due to drifts in the operating point.
@@ V Discussion
A major feature of the scheme is that it can theoretically achieve zero cross-talk between magnetic and rotation signals, although this will only hold true when the signatures are known perfectly. Achieving low cross-talk in practice means good measurement of the signatures in the first place, and then low drift of experimental parameters so the signatures do not change. A common issue in comagnetometers is drift of the compensation point due either to stray $B_{z}$ fields or drifting polarisation. For the simulation at 106.3 nT we calculate that the bias field must remain within ±$\pm 0.2\,$nT to keep the cross-talk below $0.2\,\mu$Hz/pT. This value of cross-talk corresponds to a ‘suppression’ factor of around 150 for an equivalent CW comagnetometer. This corresponds to a relative bias stability of 0.2% which is stringent, but achievable. Further, drifts in $B_{z}$ are easily detected in our scheme via the alkali precession frequency (by introducing a small transverse field if necessary), thus enabling a method for stabilising the bias field. We also expect drifts in other parameters, such as the magnetization of the spins, although the use of pulsed pumping will help in distinguishing this and other deleterious effects. It is notable that our experiment, albeit with a different mixture of gases, has already achieved a high level of magnetic suppression with only minimal optimisation. This gives us reason to be optimistic about the levels of suppression that can be achieved in future work.
@@ VI Conclusion
Through the applications of magnetic and optical pulses, we have devised a new protocol for extracting signals from an atomic comagnetometer. The scheme is able to measure magnetic fields and rotations on two orthogonal axes with a noise performance that, in ideal conditions, should be comparable to the state-of-the-art continuous wave scheme, where only a single parameter is measured. A proof-of-principle experimental realisation based on ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ was presented with promising results. In particular, we demonstrated that the PANCo scheme effectively suppresses cross-talk between magnetic and rotational signals to a level that is comparable to the best self-compensated alkali-noble comagnetometers. This result was achieved with minimal optimisation, yet the scheme has a wide parameter space for performance gains in terms of the applied pulses, the applied bias fields, gas mixtures and investigations of techniques to diagnose and minimise other technical noise. In future work we plan to explore these factors to uncover the practical limits of the scheme.

### v2

@@ I Introduction
The gyroscopic performance of the SCC has shown excellent levels of sensitivity. For example Kornack et al. [12] found a noise floor $\mathrm{5\!\times\!10^{-7}\,rad\,s^{-1}\,Hz^{-1/2}}$, which was recently surpassed by Wei et al. [23] with $\mathrm{3\!\times\!10^{-8}\,rad\,s^{-1}\,Hz^{-1/2}}$. While impressive, these results are still over an order of magnitude above the atomic shot noise limit set by the alkali, and particularly limited at low frequencies where magnetic noise is typically largest. While some progress has been made reducing magnetic and other technical noise sources [24, 25, 26, 27, 28, 29], a remaining challenge is that increased suppression of magnetic noise is not only hard to achieve, it is also difficult to maintain this suppression due to drifts in the operating point.
@@ V Discussion
Our scheme has some potential advantages over other methods with respect to gyroscopy and exotic physics searches. Rotational sensing is a key part of inertial navigation systems, but it may be augmented with magnetic navigation systems [32]. A sensor that combines both capabilities on two axes may prove useful. With respect to searches for new physics, our method has the potential to enable stronger rejection of magnetic noise, which is a key limiting factor for CW SCCs.
Working towards these applications will require careful consideration of the magnetic cross-talk. The PANCo scheme can theoretically achieve zero cross-talk between magnetic and rotation signals, although this will only hold true when the signatures are known perfectly. Achieving low cross-talk in practice means good measurement of the signatures in the first place, and then low drift of experimental parameters so the signatures do not change. A common issue in comagnetometers is drift of the compensation point due either to stray $B_{z}$ fields or drifting polarisation. For the simulation at 106.3 nT we calculate that the bias field must remain within ±$\pm 0.2\,$nT to keep the cross-talk below $0.2\,\mu$Hz/pT. This value of cross-talk corresponds to a ‘suppression’ factor of around 150 for an equivalent CW comagnetometer. This corresponds to a relative bias stability of 0.2% which is stringent, but achievable. Further, drifts in $B_{z}$ are easily detected in our scheme via the alkali precession frequency (by introducing a small transverse field if necessary), thus enabling a method for stabilising the bias field. We also expect drifts in other parameters, such as the magnetization of the spins, although the use of pulsed pumping will help in distinguishing this and other deleterious effects. It is notable that our experiment, albeit with a different mixture of gases, has already achieved a high level of magnetic suppression with only minimal optimisation. This gives us reason to be optimistic about the levels of suppression that can be achieved in future work.
@@ VI Conclusion
Through the applications of magnetic and optical pulses, we have devised a new protocol for extracting signals from an atomic comagnetometer. The scheme is able to measure magnetic and non-magnetic effects on two orthogonal axes with a noise performance that, in ideal conditions, should be comparable to the state-of-the-art continuous wave scheme, where only a single parameter is measured. A proof-of-principle experimental realisation based on ${}^{87}\mathrm{Rb}$ and ${}^{129}\mathrm{Xe}$ was presented with promising results. In particular, we demonstrated that the PANCo scheme effectively suppresses cross-talk between magnetic and non-magnetic signals to a level that is comparable to the best self-compensated alkali-noble comagnetometers. This result was achieved with minimal optimisation, yet the scheme has a wide parameter space for performance gains in terms of the applied pulses, the applied bias fields, gas mixtures and investigations of techniques to diagnose and minimise other technical noise. In future work we plan to explore these factors to uncover the practical limits of the scheme.